A 0.100 mole sample of CH4 reacts completely in a calorimeter having a heat capacity of 783 J/°C. The calorimeter contains 254 g of water. Determine the temperature increase of the calorimeter.CH4(g) + 2O2(g) -> 2H2O(g) + CO2(g) delta H° = −802.3 kJ


1
Expert's answer
2021-10-29T02:30:34-0400

∆H= m×S×∆T

Molar mass of CH4= 16.04

= 16.04×0.1= 1.604

-802.3= 0.255604×783×∆T

-802.3= 200∆T

∆T= -4.01°C